OpenStudy (anonymous):

Which expression is equivalent to 3(c + b) + 9b – c 3c + 9b 2c + 12b 2c -12b 3c – 9b

OpenStudy (anonymous):

Please help me understand this better. Cuz i'm lost U.U

OpenStudy (anonymous):

huh....

OpenStudy (anonymous):



OpenStudy (anonymous):

uh....

OpenStudy (anonymous):

Distribute

OpenStudy (anonymous):

3(c+b)= 3c+3b . So then you have to combine like terms

OpenStudy (anonymous):

3c+3b+9b-c

OpenStudy (anonymous):

3c-c=2c

OpenStudy (anonymous):

3b+9b=12b

OpenStudy (anonymous):

oh ok. So the answer is 3b+9b=12b...?

OpenStudy (anonymous):

No. The answer is 2c+12b

OpenStudy (anonymous):

really O.O